# On the Hausdorff dimension and Cantor set structure of sliding Shilnikov invariant sets

### v2

@@ 3.2. The First Return Map as a CIFS
There exists a neighborhood $U\subseteq\gamma_{r}$ of $q$ satisfying (P) for which the $\Psi_{U}$ is conformal.
@@
In what follows, we will examine the conformal conditions individually.
Let $U\subseteq\gamma_{r}$ be the neighborhood of $q$ given by Proposition F.
@@
Consider the set $\gamma_{r+\delta}$ for $\delta>0$, and for each $J\in\cal{J}_{U}$, define the extension $\psi_{J}^{\delta}:\gamma_{r+\delta}\to\gamma_{r+\delta}$ of $\psi_{J}$ by joining affine functions with slopes equal to the lateral derivatives of $\psi_{J}$ at the endpoints of $\gamma_{r}$ on each side. This extension is a diffeomorphims, and since $\gamma_{r+\delta}$ is compact, the derivative $(\psi_{J}^{\delta})^{\prime}$ is Lipschitz continuous and, consequently, $\psi_{J}^{\delta}$ is also $\cal{C}^{1,\varepsilon}$. Furthermore, by Proposition F, $(\psi_{J}^{\delta})^{\prime}(x)\neq 0$ for every $x\in\gamma_{r}$, and therefore, for every $x\in\gamma_{r+\delta}$. Finally, since we are dealing with functions defined on a one-dimensional set, such extensions are trivially conformal, satisfying the requirements for Condition (C4) to hold.
At last, we will show that Condition (C6) holds. Notice that, for $x,y\in\gamma_{r}$, we have
@@
$\psi^{\prime}_{J}(x)=\Big{[}\left(\rho_{X}^{-1}\right)^{\prime}(f_{G_{X}}(x))\Big{]}\cdot\Big{[}\left(H^{-1}\right)^{\prime}(f_{H}(x))\Big{]}\cdot\lambda^{-c_{J}}\cdot\Big{[}H^{\prime}\left(\left(\rho^{J}_{\tilde{Z}}\right)^{-1}(x)\right)\Big{]}\cdot\Big{[}\left(\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)^{\prime}(x)\Big{]},$
where $f_{G_{X}}:=H^{-1}\left(\lambda^{-c_{J}}H\circ\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)$ and $f_{H}:=\lambda^{-c_{J}}H\circ\left(\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)$. This implies that
$\inf_{x\in\gamma_{r}}\left\lvert{\psi^{\prime}_{J}(x)}\right\rvert\geq\lambda^{-c_{J}}\cdot C,$
where $C:=\left(\inf_{\zeta\in\mu_{r}}\Big{\lvert}\left(\rho_{X}^{-1}\right)^{\prime}(\zeta)\Big{\rvert}\right)\cdot\left(\inf_{z\in H(\mu_{r})}\Big{\lvert}(H^{-1})^{\prime}(z)\Big{\rvert}\right)\cdot\left(\inf_{\zeta\in\mu_{r}}\Big{\lvert}H^{\prime}(\zeta)\Big{\rvert}\right)\cdot\left(\inf_{x\in\gamma_{r}}\Big{\lvert}\left(G^{J}_{\tilde{Z}}\right)^{\prime}(x)\Big{\rvert}\right)>0$, given that every function in the composition is a diffeomorphism.
Let us denote $T(x):=\lambda^{c_{J}}\cdot\psi_{J}(x)$. Notice that $T(x)$ is $\cal{C}^{1,\alpha}$, for some $\alpha>0$, since each of the functions $\rho_{\tilde{Z}}^{-1}$ and $\rho_{X}^{-1}$ is a $\cal{C}^{1,1}$ diffeomorphism (the same class of differentiability of the fields $X,Y$ that induce them), and $H,H^{-1}$ are $\cal{C}^{1,\beta}$ maps, and in compact domains, products and compositions preserves the Hölder condition, with some adjustments in the multiplicative and exponential constants (this can be proven using many times the Lipschitz and Hölder conditions). Keeping this in mind, we deduce the following inequality:
@@
for some $D>1,\alpha>0$. We can, then, extend the inequality for $\psi_{J}^{\delta}$ and $x,y\in\gamma_{r+\delta}$, since the derivatives $(\psi_{J}^{\delta})^{\prime}$ in $\gamma_{r+\delta}$ attain values already obtained by $\psi_{J}^{\prime}$ in $\gamma_{r}$ (by construction), so we finally deduce (C6).
Therefore, our IFS satisfies all the necessary conditions, making it CIFS. es all the necessary conditions, making it CIFS. ∎
Now, we will demonstrate that the attractor set of CIFS $\Psi_{U}$, given by (5), coincides with the invariant set of the restricted first return map $\pi_{U}:=\pi|_{\cal{W}_{r}\cap U}$ given by (3).
@@
The invariant set $\Lambda_{U}$ of $\pi_{U}$ coincide with the attractor set $\Delta_{U}$ of the CIFS $\Psi_{U}$.
@@ 3.3. The Closure of $\Lambda_{U}$
$\Delta_{U}^{k+1}=\bigcup_{\eta\in\cal{J}_{U}^{k+1}}\psi_{\eta}(\gamma_{r})=\bigcup_{J\in\cal{J}_{U}}\psi_{J}(\Delta_{U}^{k}),\quad\text{for any $k\geq 1$},$ (15)
In order provide a characterization for the closure $\overline{\Delta_{U}}$, define
@@
The sets $\Delta_{U}^{k}\cup Q_{U}^{k-1}$, for $k\geq 1$, and $\Lambda_{U}^{q}$ are compact.
@@
First, notice that the relationships 15 and 17 imply
@@
Since $\psi(x_{i})\in\Delta_{U}$, because $\Delta_{U}$ is $\pi$-invariant, we conclude that $x$ the limit of a sequence of elements of $\Delta_{U}$, therefore $x\in\overline{\Delta_{U}}$. ∎
@@
$Q_{U}=\left(\bigcup_{k\geq 0}\pi^{-k}(q)\right)\cap U=\bigcup_{k\geq 0}\pi_{U}^{-k}(q).$ (18)
@@ 3.4. Proof of Theorem A
which, taking (21) into account, implies $P_{U}(1)<1$. Thus, there exists $10<\tilde{t}<$ 1 such that $P(\tilde{t})=1$.
@@
Clearly, $\overline{\Lambda_{U}}$ inherits the metric structure from $\gamma_{r}$ and is compact. Besides, given that $\operatorname{dim_{H}}\left(\overline{\Lambda_{U}}\right)<1$, Property (B5) says that $\overline{\Lambda_{U}}$ is a totally disconnected set. In order to see that it is perfect, let $x\in\Lambda_{U}$. From (5), we know that $x$ is represented by some $\overline{\eta}=(J_{1},J_{2},J_{3},\ldots)\in\cal{J}_{U}^{\mathbb{N}}$, that is, $x=p(\overline{\eta})$, where $p$ is the projection defined in (4). Taking a sequence $\left\{x_{j}\right\}_{j\geq 1}\subseteq\Lambda_{U}$ satisfying $x_{k}\in\psi_{(J_{1},\ldots,J_{k})}(\gamma_{r})$, but $x_{k}\not\in\psi_{(J_{1},\ldots,J_{k},J_{k+1})}(\gamma_{r})$, for $k\geq 1$, we have that $\left\{x_{j}\right\}_{j\geq 1}\subseteq\Lambda_{U}\setminus\left\{x\right\}$ and $x_{j}\to x$. This implies that every point of $\Lambda_{U}$ is an accumulation point. Since the closure of a set does not add any isolated points, $\overline{\Lambda_{U}}$ is a perfect set. This concludes the proof of Theorem. ∎

### v4

@@ 3.2. The First Return Map as a CIFS
There exists a neighborhood $U\subseteq\gamma_{r}$ of $q$ satisfying (P) for which IFS $\Psi_{U}$ is conformal.
@@
In what follows, we will examine the conformal conditions individually. Let $U\subseteq\gamma_{r}$ be the neighborhood of $q$ given by Proposition F.
@@
Consider the set $\gamma_{r+\delta}$ for $\delta>0$, and for each $J\in\cal{J}_{U}$, define the extension $\psi_{J}^{\delta}:\gamma_{r+\delta}\to\gamma_{r+\delta}$ of $\psi_{J}$ by joining affine functions with slopes equal to the lateral derivatives of $\psi_{J}$ at the endpoints of $\gamma_{r}$ on each side. This extension is a diffeomorphism, and since $\gamma_{r+\delta}$ is compact, the derivative $(\psi_{J}^{\delta})^{\prime}$ is Lipschitz continuous and, consequently, $\psi_{J}^{\delta}$ is also $\cal{C}^{1,\varepsilon}$. Furthermore, by Proposition F, $(\psi_{J}^{\delta})^{\prime}(x)\neq 0$ for every $x\in\gamma_{r}$, and therefore, for every $x\in\gamma_{r+\delta}$. Finally, since we are dealing with functions defined on a one-dimensional set, such extensions are trivially conformal, satisfying the requirements for Condition (C4) to hold.
At last, we will show that Condition (C6) holds. Notice that we have
@@
$\psi^{\prime}_{J}(x)=\Big{[}\left(\rho_{X}^{-1}\right)^{\prime}(f_{H^{-1}}(x))\Big{]}\cdot\Big{[}\left(H^{-1}\right)^{\prime}(f_{H}(x))\Big{]}\cdot\lambda^{-c_{J}}\cdot\Big{[}H^{\prime}\left(\left(\rho^{J}_{\tilde{Z}}\right)^{-1}(x)\right)\Big{]}\cdot\Big{[}\left(\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)^{\prime}(x)\Big{]},$
where $f_{H^{-1}}:=H^{-1}\left(\lambda^{-c_{J}}H\circ\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)$ and $f_{H}:=\lambda^{-c_{J}}H\circ\left(\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)$.
Similarly to (12), given that each function in the composition is a diffeomorphism, we can set
$C:=\bigg{(}\min_{\zeta\in\mu_{r}}\left\lvert{\left(\rho_{X}^{-1}\right)^{\prime}(\zeta)}\right\rvert\bigg{)}\cdot\bigg{(}\min_{z\in H(\mu_{r})}\left\lvert{(H^{-1})^{\prime}(z)}\right\rvert\bigg{)}\cdot\bigg{(}\min_{\zeta\in\mu_{r}}\left\lvert{H^{\prime}(\zeta)}\right\rvert\bigg{)}\cdot\bigg{(}\min_{x\in\gamma_{r}}\left\lvert{\left(\left(\rho_{\tilde{Z}}^{J}\right)^{-1}\right)^{\prime}(x)}\right\rvert\bigg{)}>0.$
This implies that
$\inf_{x\in\gamma_{r}}\left\lvert{\psi^{\prime}_{J}(x)}\right\rvert\geq C\,\lambda^{-c_{J}}.$ (14)
Let us denote $T(x):=\lambda^{c_{J}}\cdot\psi_{J}(x)$. Notice that $T(x)$ is $\cal{C}^{1,\alpha}$, for some $\alpha>0$, since each of the functions $\rho_{\tilde{Z}}^{-1}$ and $\rho_{X}^{-1}$ is a $\cal{C}^{1,1}$ diffeomorphism (the same class of differentiability of the fields $X$ and $Y$ that induce them), and $H,H^{-1}$ are $\cal{C}^{1,\beta}$ maps, and in compact domains, products and compositions preserve the Hölder condition, with some adjustments in the multiplicative and exponential constants (this can be proven using many times the Lipschitz and Hölder conditions). Keeping this in mind and applying in (14), we deduce the following inequality:
@@
for some $D>1,\alpha>0$ and for all $x,y\in\gamma_{r}$,. We can, then, extend the inequality for $\psi_{J}^{\delta}$ and $x,y\in\gamma_{r+\delta}$, since the derivatives $(\psi_{J}^{\delta})^{\prime}$ in $\gamma_{r+\delta}$ attain values already obtained by $\psi_{J}^{\prime}$ in $\gamma_{r}$ (by construction), so we finally deduce (C6).
Therefore, our IFS satisfies all the necessary conditions, making it a CIFS. ∎
Now, we will demonstrate that the attractor set of the CIFS $\Psi_{U}$, given by (6), coincides with the invariant set of the restricted first return map $\pi_{U}:=\pi|_{\cal{W}_{r}\cap U}$ given by (4).
@@
The invariant set $\Lambda_{U}$ of $\pi_{U}$ coincides with the attractor set $\Delta_{U}$ of the CIFS $\Psi_{U}$.
@@ 3.3. The Closure of $\Lambda_{U}$
$\Delta_{U}^{k+1}=\bigcup_{\eta\in\cal{J}_{U}^{k+1}}\psi_{\eta}(\gamma_{r})=\bigcup_{J\in\cal{J}_{U}}\psi_{J}(\Delta_{U}^{k}),\quad\text{for any $k\geq 1$}.$ (18)
In order to provide a characterization for the closure $\overline{\Delta_{U}}$, define
@@
The sets $\Delta_{U}^{k}\cup Q_{U}^{k-1}$, for $k\geq 1$, and $\Delta_{U}^{q}$ are compact.
@@
First, notice that the relationships (18) and (20) imply
@@
Since $\psi_{\overline{\eta}}(x_{i})\in\Delta_{U}$, because $\Delta_{U}$ is $\pi$-invariant, we conclude that $x$ the limit of a sequence of elements of $\Delta_{U}$, therefore $x\in\overline{\Delta_{U}}$. ∎
@@
$Q_{U}=\left(\bigcup_{k\geq 0}\pi^{-k}(q)\right)\cap U=\bigcup_{k\geq 0}\pi_{U}^{-k}(q),$ (21)
@@ 3.4. Proof of Theorem A
which, taking (24) into account, implies $P_{U}(1)<1$. Thus, there exists $10<\tilde{t}<$ 1 such that $P_{U}(\tilde{t})=1$.
@@
Clearly, $\overline{\Lambda_{U}}$ inherits the metric structure from $\gamma_{r}$ and is compact. Besides, given that $\operatorname{dim_{H}}\left(\overline{\Lambda_{U}}\right)<1$, Property (B5) says that $\overline{\Lambda_{U}}$ is a totally disconnected set. In order to see that it is perfect, let $x\in\Lambda_{U}$. From (6), we know that $x$ is represented by some $\overline{\eta}=(J_{1},J_{2},J_{3},\ldots)\in\cal{J}_{U}^{\mathbb{N}}$, that is, $x=\operatorname{proj}(\overline{\eta})$, where $\operatorname{proj}$ denotes the projection defined in (5), which is well-defined due to the uniformly contractive property (C2) satisfied by the CIFS $\Psi_{U}$ (see Remark 2). Taking a sequence $\left\{x_{j}\right\}_{j\geq 1}\subseteq\Lambda_{U}$ satisfying $x_{k}\in\psi_{(J_{1},\ldots,J_{k})}(\gamma_{r})$, but $x_{k}\not\in\psi_{(J_{1},\ldots,J_{k},J_{k+1})}(\gamma_{r})$, for $k\geq 1$, we have that $\left\{x_{j}\right\}_{j\geq 1}\subseteq\Lambda_{U}\setminus\left\{x\right\}$ and $x_{j}\to x$. This implies that every point of $\Lambda_{U}$ is an accumulation point. Since the closure of a set does not add any isolated points, $\overline{\Lambda_{U}}$ is a perfect set. This concludes the proof. ∎